OpenStudy (anonymous):

A 2 mi. cab ride costs $ 5.25. A 5 mi. cab ride costs $10.50. which equation models the cost y of the cab ride for ride that is x miles? a. y = 1.75x - 1.75 b. y = 1.75x + 1.75 c. y = -1.75x - 1.75 d. y = -1.75x + 1.75

OpenStudy (anonymous):

@Starrhazel1234 please help will reward and fan

OpenStudy (anonymous):

I say D @jimmiedeel

OpenStudy (anonymous):

I think the trick here is to realize that this cost and mileage are forming two points on a line so while it's not flat rate per dollar there is a gradual increasing cost (this would also make sense as when you get a taxi there is usually a fee you pay just to get in and the mileage builds from there).

OpenStudy (anonymous):

YOU A DAB

OpenStudy (anonymous):

What I did to solve this is find the equation of the line using slope intercept form y is the cost of the ride which increases as you increase mileage or move along the X axis. My points are (2, 5.25) and (5,10.5) y=mx+b m=y2-y1/x2-x1 m=(10.5-5.25)/(5-2) m=5.25/3=1.75 or 7/4 y=7/4x+b enter one point for the values of x and y to find b 10.5=(1.75X5) + b 10.5=8.75 + b b=1.75 The equation of the line is y=7/4x+1.75 (I also double checked using the second point) Then all you have to do is enter the mileage 3.8 for the value of X and solve for y y=(7/4 X 3.8) + 1.75 y=6.65+1.75 y=8.4 The cab ride for the 3.8mile trip is $8.4 Does this help @jimmiedeel

OpenStudy (anonymous):

Do I get a medal @jimmiedeel ?

OpenStudy (anonymous):

No b/c that was not the right answer or solution